OpenStudy (anonymous):

Find the maximum value of 3 – 2 cos X

OpenStudy (solomonzelman):

This is not the only way, if you explain the question, and then the asker tells you "but what's the answer?" or something along this line, then you know ....

OpenStudy (solomonzelman):

Anyway.... if I were to deal with a triangle it would be less than five....

OpenStudy (anonymous):

how cosx =1 and 2cosx =2

OpenStudy (solomonzelman):

I mean cos(X)=-1 2cos(X)= -2 3-(-2) = 3+2=5

OpenStudy (anonymous):

:find the absolute min and max values of f(x)= 2Cosx + Sin2x on the interval [0,pi/2]? Solution : f(x)= 2Cosx + Sin2x =>f'(x)=d/dx( 2Cosx + Sin2x) =-2sinx+2cos2x now,f'(x)=0 => -2sinx+2cos2x=0 =>-2sinx+2(1-2sin^2x)=0 [cos2x= 1-2sin^2x] => -2sinx+2-4sin^2x+2=0 => 4sin^2x+2sinx-2=0 =>4sin^2x+4sinx-2sinx-2=0 => 4sinx(sinx+1)-2(sinx+1)=0 =>(4sinx-2)(sinx+1)=0 => 4sinx=2 or sinx=-1 [condition not applicable for -ve ] =>sinx=1/2=>x=π/6 Now, f(0)=2cos(0)+sin(2*0) =2*1+0 =2 f(π/6)= 2cos(π/6)+sin(2*π/6) = 2.√3/2+√3/2 = √3+√3/2 =( 2√3+√3)/2 =3√3/2 f(π/2)= 2cos(π/2)+sin(2*π/2) = 2*0+0=0 Maximum value 3√3/2,at x=π/6 Minimum value 0 ,at x=π/2

OpenStudy (anonymous):

kk

OpenStudy (solomonzelman):

No, it is shifted 3 units up .

OpenStudy (solomonzelman):

yeah, don't leave that out :)

OpenStudy (solomonzelman):

I am going to post some shift rules on a different example. \(\large\color{ blue }{\large {\bbox[5pt, lightyellow ,border:2px solid white ]{ \large\text{ }\\ \begin{array}{|c|c|c|c|} \hline~~~~~~~~~~~~~~~~~~~~~~~~~~~\textbf{Shifts}~~~~~~~~~~~~~~~~~~~~~~~~~~~&~\bf{c~~~units~~~~} \\ \hline \\f(x)= ∜x ~~~ ⇒ ~~~ f(x)= \sqrt[4]{x \normalsize\color{red }{ -~\rm{c}} } &~\rm{to~~the~~right~} \\ \text{ } \\ f(x)= ∜x ~~~ ⇒ ~~~ f(x)= \sqrt[4]{x \normalsize\color{red}{ +~\rm{c}} } &~\rm{to~~the~~left ~} \\ \text{ } \\ f(x)= ∜x ~~~ ⇒ ~~~ f(x)= ∜x \normalsize\color{red}{ +~\rm{c} } &~\rm{up~} \\ \text{ } \\ f(x)= ∜x ~~~ ⇒ ~~~ f(x)= ∜x \normalsize\color{red}{ -~\rm{c} } &~\rm{down~} \\ \\ \hline \end{array} }}}\)

OpenStudy (solomonzelman):

no need for desmos, but ty

OpenStudy (solomonzelman):

my favorite graphing tool (after my paper and TI-84 calc)

OpenStudy (solomonzelman):

oH yeah

OpenStudy (solomonzelman):

I sometimes graph in the question if I have time....

OpenStudy (anonymous):

idk shift rules and how to use it.and idk about graph

OpenStudy (anonymous):

explain @SolomonZelman

OpenStudy (anonymous):

wowee, this is not nearly so complicated i don't think

OpenStudy (anonymous):

Find the maximum value of \(3 - 2 \cos(x)\) the largest cosine can be is 1 and the smallest is -1 the biggest \(-2\cos(x)\) can be is \(2\) the biggest \(3-2\cos(x)\) can be is \(3+2=5\)

OpenStudy (solomonzelman):

yes I said that already, I was just giving an example of shifts ...